\begin{thebibliography}{67}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bai et~al.(2023)Bai, Wang, Yin, Sun, Chen, Li, and Li]{bai2023deep}
Ting Bai, Le~Wang, Dameng Yin, Kaimin Sun, Yepei Chen, Wenzhuo Li, and Deren Li.
\newblock Deep learning for change detection in remote sensing: a review.
\newblock \emph{Geo-spatial Information Science}, 26\penalty0 (3):\penalty0 262--288, 2023.

\bibitem[Bandara(2023)]{wgcbanchangeformergithubissue1}
Chaminda Bandara.
\newblock Changeformer\_cd source code data leakage issue 1.
\newblock \url{https://github.com/wgcban/ChangeFormer/issues/52}, 2023.

\bibitem[Bandara \& Patel(2022{\natexlab{a}})Bandara and Patel]{bandara2022revisiting}
Wele Gedara~Chaminda Bandara and Vishal~M Patel.
\newblock Revisiting consistency regularization for semi-supervised change detection in remote sensing images.
\newblock \emph{arXiv preprint arXiv:2204.08454}, 2022{\natexlab{a}}.

\bibitem[Bandara \& Patel(2022{\natexlab{b}})Bandara and Patel]{bandara2022transformer}
Wele Gedara~Chaminda Bandara and Vishal~M Patel.
\newblock A transformer-based siamese network for change detection.
\newblock In \emph{IGARSS 2022-2022 IEEE International Geoscience and Remote Sensing Symposium}, pp.\  207--210. IEEE, 2022{\natexlab{b}}.

\bibitem[Bello et~al.(2021)Bello, Fedus, Du, Cubuk, Srinivas, Lin, Shlens, and Zoph]{bello2021revisiting}
Irwan Bello, William Fedus, Xianzhi Du, Ekin~Dogus Cubuk, Aravind Srinivas, Tsung-Yi Lin, Jonathon Shlens, and Barret Zoph.
\newblock Revisiting resnets: Improved training and scaling strategies.
\newblock \emph{Advances in Neural Information Processing Systems}, 34:\penalty0 22614--22627, 2021.

\bibitem[Cambrin et~al.(2024)Cambrin, Colomba, and Garza]{cambrin2024cabuar}
Daniele~Rege Cambrin, Luca Colomba, and Paolo Garza.
\newblock Cabuar: California burned areas dataset for delineation.
\newblock \emph{arXiv preprint arXiv:2401.11519}, 2024.

\bibitem[Chen \& Shi(2020)Chen and Shi]{chen2020spatial}
Hao Chen and Zhenwei Shi.
\newblock A spatial-temporal attention-based method and a new dataset for remote sensing image change detection.
\newblock \emph{Remote Sensing}, 12\penalty0 (10):\penalty0 1662, 2020.

\bibitem[Chen et~al.(2021{\natexlab{a}})Chen, Li, and Shi]{chen2021adversarial}
Hao Chen, Wenyuan Li, and Zhenwei Shi.
\newblock Adversarial instance augmentation for building change detection in remote sensing images.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 60:\penalty0 1--16, 2021{\natexlab{a}}.

\bibitem[Chen et~al.(2021{\natexlab{b}})Chen, Qi, and Shi]{chen2021remote}
Hao Chen, Zipeng Qi, and Zhenwei Shi.
\newblock Remote sensing image change detection with transformers.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 60:\penalty0 1--14, 2021{\natexlab{b}}.

\bibitem[Chen et~al.(2023)Chen, Zhang, Chen, Zhou, Chen, Zhou, and Shi]{chen2023remote}
Hao Chen, Haotian Zhang, Keyan Chen, Chenyao Zhou, Song Chen, Zhengxia Zhou, and Zhenwei Shi.
\newblock Remote sensing image change detection towards continuous bitemporal resolution differences.
\newblock \emph{arXiv preprint arXiv:2305.14722}, 2023.

\bibitem[Codegoni et~al.(2023)Codegoni, Lombardi, and Ferrari]{codegoni2023tinycd}
Andrea Codegoni, Gabriele Lombardi, and Alessandro Ferrari.
\newblock Tinycd: a (not so) deep learning model for change detection.
\newblock \emph{Neural Computing and Applications}, 35\penalty0 (11):\penalty0 8471--8486, 2023.

\bibitem[Daudt et~al.(2018)Daudt, Le~Saux, and Boulch]{daudt2018fully}
Rodrigo~Caye Daudt, Bertr Le~Saux, and Alexandre Boulch.
\newblock Fully convolutional siamese networks for change detection.
\newblock In \emph{2018 25th IEEE International Conference on Image Processing (ICIP)}, pp.\  4063--4067. IEEE, 2018.

\bibitem[Deng et~al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei]{deng2009imagenet}
Jia Deng, Wei Dong, Richard Socher, Li-Jia Li, Kai Li, and Li~Fei-Fei.
\newblock Imagenet: A large-scale hierarchical image database.
\newblock In \emph{2009 IEEE conference on computer vision and pattern recognition}, pp.\  248--255. Ieee, 2009.

\bibitem[Du et~al.(2021)Du, Li, Cui, Qian, Li, and Bello]{du2021revisiting}
Xianzhi Du, Yeqing Li, Yin Cui, Rui Qian, Jing Li, and Irwan Bello.
\newblock Revisiting 3d resnets for video recognition.
\newblock \emph{arXiv preprint arXiv:2109.01696}, 2021.

\bibitem[Fang et~al.(2023)Fang, Li, and Li]{fang2023changer}
Sheng Fang, Kaiyu Li, and Zhe Li.
\newblock Changer: Feature interaction is what you need for change detection.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 2023.

\bibitem[Fazry et~al.(2023)Fazry, Ramadhan, and Jatmiko]{fazry2023change}
Lhuqita Fazry, Mgs M~Luthfi Ramadhan, and Wisnu Jatmiko.
\newblock Change detection of high-resolution remote sensing images through adaptive focal modulation on hierarchical feature maps.
\newblock \emph{IEEE Access}, 2023.

\bibitem[Gerard et~al.(2024)Gerard, Borne-Pons, and Sullivan]{gerard2024simple}
Sebastian Gerard, Paul Borne-Pons, and Josephine Sullivan.
\newblock A simple, strong baseline for building damage detection on the xbd dataset.
\newblock \emph{arXiv preprint arXiv:2401.17271}, 2024.

\bibitem[Gupta et~al.(2019{\natexlab{a}})Gupta, Goodman, Patel, Hosfelt, Sajeev, Heim, Doshi, Lucas, Choset, and Gaston]{gupta2019creating}
Ritwik Gupta, Bryce Goodman, Nirav Patel, Ricky Hosfelt, Sandra Sajeev, Eric Heim, Jigar Doshi, Keane Lucas, Howie Choset, and Matthew Gaston.
\newblock Creating xbd: A dataset for assessing building damage from satellite imagery.
\newblock In \emph{Proceedings of the IEEE/CVF conference on computer vision and pattern recognition workshops}, pp.\  10--17, 2019{\natexlab{a}}.

\bibitem[Gupta et~al.(2019{\natexlab{b}})Gupta, Hosfelt, Sajeev, Patel, Goodman, Doshi, Heim, Choset, and Gaston]{gupta2019xbd}
Ritwik Gupta, Richard Hosfelt, Sandra Sajeev, Nirav Patel, Bryce Goodman, Jigar Doshi, Eric Heim, Howie Choset, and Matthew Gaston.
\newblock xbd: A dataset for assessing building damage from satellite imagery.
\newblock \emph{arXiv preprint arXiv:1911.09296}, 2019{\natexlab{b}}.

\bibitem[He et~al.(2016)He, Zhang, Ren, and Sun]{he2016deep}
Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun.
\newblock Deep residual learning for image recognition.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and pattern recognition}, pp.\  770--778, 2016.

\bibitem[Henderson et~al.(2018)Henderson, Islam, Bachman, Pineau, Precup, and Meger]{henderson2018deep}
Peter Henderson, Riashat Islam, Philip Bachman, Joelle Pineau, Doina Precup, and David Meger.
\newblock Deep reinforcement learning that matters.
\newblock In \emph{Proceedings of the AAAI conference on artificial intelligence}, volume~32, 2018.

\bibitem[Huang et~al.(2024)Huang, Xu, and Zhang]{huang2024remote}
Bin Huang, Yichen Xu, and Feng Zhang.
\newblock Remote sensing image change detection based on adjacent-level feature fusion and dense skip connections.
\newblock \emph{IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing}, 2024.

\bibitem[Ji et~al.(2018)Ji, Wei, and Lu]{ji2018fully}
Shunping Ji, Shiqing Wei, and Meng Lu.
\newblock Fully convolutional networks for multisource building extraction from an open aerial and satellite imagery data set.
\newblock \emph{IEEE Transactions on geoscience and remote sensing}, 57\penalty0 (1):\penalty0 574--586, 2018.

\bibitem[Koch et~al.(2015)Koch, Zemel, Salakhutdinov, et~al.]{koch2015siamese}
Gregory Koch, Richard Zemel, Ruslan Salakhutdinov, et~al.
\newblock Siamese neural networks for one-shot image recognition.
\newblock In \emph{ICML deep learning workshop}, volume~2. Lille, 2015.

\bibitem[Lacoste et~al.(2023)Lacoste, Lehmann, Rodriguez, Sherwin, Kerner, L{\"u}tjens, Irvin, Dao, Alemohammad, Drouin, et~al.]{lacoste2023geo}
Alexandre Lacoste, Nils Lehmann, Pau Rodriguez, Evan~David Sherwin, Hannah Kerner, Bj{\"o}rn L{\"u}tjens, Jeremy~Andrew Irvin, David Dao, Hamed Alemohammad, Alexandre Drouin, et~al.
\newblock Geo-bench: Toward foundation models for earth monitoring.
\newblock \emph{arXiv preprint arXiv:2306.03831}, 2023.

\bibitem[Li \& Wu()Li and Wu]{li4779358using}
Jialu Li and Chen Wu.
\newblock Using difference features effectively: A multi-task network for exploring change areas and change moments in time series remote sensing images.
\newblock \emph{Available at SSRN 4779358}.

\bibitem[Li(2022)]{opencd}
Kaiyu Li.
\newblock Opencd.
\newblock \url{https://github.com/likyoo/open-cd}, 2022.

\bibitem[Li et~al.(2024)Li, Wang, Fang, Zhao, Yang, and Li]{li2024decoder}
Zhe Li, Xiaoxin Wang, Sheng Fang, Jianli Zhao, Shuqi Yang, and Wen Li.
\newblock A decoder-focused multi-task network for semantic change detection.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 2024.

\bibitem[Li et~al.(2023)Li, Tang, Liu, Li, Li, and Zhang]{li2023ms}
Zhenglai Li, Chang Tang, Xinwang Liu, Changdong Li, Xianju Li, and Wei Zhang.
\newblock Ms-former: Memory-supported transformer for weakly supervised change detection with patch-level annotations.
\newblock \emph{arXiv preprint arXiv:2311.09726}, 2023.

\bibitem[Lin et~al.(2024)Lin, Hang, Wang, and Liu]{lin2024diformer}
Hui Lin, Renlong Hang, Shanmin Wang, and Qingshan Liu.
\newblock Diformer: A difference transformer network for remote sensing change detection.
\newblock \emph{IEEE Geoscience and Remote Sensing Letters}, 2024.

\bibitem[Liu et~al.(2023{\natexlab{a}})Liu, Lin, Liu, Yu, and Li]{liu2023attention}
Wei Liu, Yiyuan Lin, Weijia Liu, Yongtao Yu, and Jonathan Li.
\newblock An attention-based multiscale transformer network for remote sensing image change detection.
\newblock \emph{ISPRS Journal of Photogrammetry and Remote Sensing}, 202:\penalty0 599--609, 2023{\natexlab{a}}.

\bibitem[Liu et~al.(2023{\natexlab{b}})Liu, Liu, Jiao, Li, Liu, and Zhang]{liu2023swin}
Xu~Liu, Yu~Liu, Licheng Jiao, Lingling Li, Fang Liu, and Dan Zhang.
\newblock Swin resnetswin transformers for change detection in remote sensing images.
\newblock In \emph{IGARSS 2023-2023 IEEE International Geoscience and Remote Sensing Symposium}, pp.\  6660--6663. IEEE, 2023{\natexlab{b}}.

\bibitem[Liu et~al.(2024)Liu, Liu, Jiao, Li, Liu, Yang, and Hou]{liu2024mutsimnet}
Xu~Liu, Yu~Liu, Licheng Jiao, Lingling Li, Fang Liu, Shuyuan Yang, and Biao Hou.
\newblock Mutsimnet: Mutually reinforcing similarity learning for rs image change detection.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 2024.

\bibitem[Liu et~al.(2020)Liu, Pang, Zhan, Zhang, and Yang]{liu2020building}
Yi~Liu, Chao Pang, Zongqian Zhan, Xiaomeng Zhang, and Xue Yang.
\newblock Building change detection for remote sensing images using a dual-task constrained deep siamese convolutional network model.
\newblock \emph{IEEE Geoscience and Remote Sensing Letters}, 18\penalty0 (5):\penalty0 811--815, 2020.

\bibitem[Liu et~al.(2023{\natexlab{c}})Liu, Li, Xiao, Huang, and Yang]{liu2023feature}
Yikun Liu, Mingsong Li, Tao Xiao, Yuwen Huang, and Gongping Yang.
\newblock Feature alignment and refinement for remote sensing images change detection.
\newblock \emph{International Journal of Remote Sensing}, 44\penalty0 (24):\penalty0 7827--7856, 2023{\natexlab{c}}.

\bibitem[Lu et~al.(2024)Lu, Wei, and Nguyen]{lu2024bi}
Wen Lu, Lu~Wei, and Minh Nguyen.
\newblock Bi-temporal attention transformer for building change detection and building damage assessment.
\newblock \emph{IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing}, 2024.

\bibitem[Musgrave et~al.(2020)Musgrave, Belongie, and Lim]{musgrave2020metric}
Kevin Musgrave, Serge Belongie, and Ser-Nam Lim.
\newblock A metric learning reality check.
\newblock In \emph{Computer Vision--ECCV 2020: 16th European Conference, Glasgow, UK, August 23--28, 2020, Proceedings, Part XXV 16}, pp.\  681--699. Springer, 2020.

\bibitem[Musgrave et~al.(2021)Musgrave, Belongie, and Lim]{musgrave2021unsupervised}
Kevin Musgrave, Serge Belongie, and Ser-Nam Lim.
\newblock Unsupervised domain adaptation: A reality check.
\newblock \emph{arXiv preprint arXiv:2111.15672}, 2021.

\bibitem[Pang et~al.(2024)Pang, Weng, Wu, Wang, and Xia]{pang2024hicd}
Chao Pang, Xingxing Weng, Jiang Wu, Qiang Wang, and Gui-Song Xia.
\newblock Hicd: Change detection in quality-varied images via hierarchical correlation distillation.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 62:\penalty0 1--16, 2024.

\bibitem[Peng et~al.(2023)Peng, Shi, Zhang, and Wang]{peng2023fda}
Wenguang Peng, Wenzhong Shi, Min Zhang, and Lukang Wang.
\newblock Fda-ffnet: A feature-distance attention based change detection network for remote sensing image.
\newblock \emph{IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing}, 2023.

\bibitem[Quan et~al.(2023)Quan, Yu, Guo, Lu, Jiang, Zheng, and He]{quan2023unified}
Yujun Quan, Anzhu Yu, Wenyue Guo, Xuanbei Lu, Bingchun Jiang, Shulei Zheng, and Peipei He.
\newblock Unified building change detection pre-training method with masked semantic annotations.
\newblock \emph{International Journal of Applied Earth Observation and Geoinformation}, 120:\penalty0 103346, 2023.

\bibitem[Ren et~al.(2024)Ren, Wang, Xia, and Lin]{ren2024mfinet}
Wuxu Ren, Zhongchen Wang, Min Xia, and Haifeng Lin.
\newblock Mfinet: Multi-scale feature interaction network for change detection of high-resolution remote sensing images.
\newblock \emph{Remote Sensing}, 16\penalty0 (7):\penalty0 1269, 2024.

\bibitem[Ronneberger et~al.(2015)Ronneberger, Fischer, and Brox]{ronneberger2015u}
Olaf Ronneberger, Philipp Fischer, and Thomas Brox.
\newblock U-net: Convolutional networks for biomedical image segmentation.
\newblock In \emph{Medical Image Computing and Computer-Assisted Intervention--MICCAI 2015: 18th International Conference, Munich, Germany, October 5-9, 2015, Proceedings, Part III 18}, pp.\  234--241. Springer, 2015.

\bibitem[Shen et~al.(2021)Shen, Lu, Chen, Wei, Xie, Yue, Chen, Lv, and Jiang]{shen2021s2looking}
Li~Shen, Yao Lu, Hao Chen, Hao Wei, Donghai Xie, Jiabao Yue, Rui Chen, Shouye Lv, and Bitao Jiang.
\newblock S2looking: A satellite side-looking dataset for building change detection.
\newblock \emph{Remote Sensing}, 13\penalty0 (24):\penalty0 5094, 2021.

\bibitem[Srivastava et~al.(2022)Srivastava, Rastogi, Rao, Shoeb, Abid, Fisch, Brown, Santoro, Gupta, Garriga-Alonso, et~al.]{srivastava2022beyond}
Aarohi Srivastava, Abhinav Rastogi, Abhishek Rao, Abu Awal~Md Shoeb, Abubakar Abid, Adam Fisch, Adam~R Brown, Adam Santoro, Aditya Gupta, Adri{\`a} Garriga-Alonso, et~al.
\newblock Beyond the imitation game: Quantifying and extrapolating the capabilities of language models.
\newblock \emph{arXiv preprint arXiv:2206.04615}, 2022.

\bibitem[Stewart et~al.(2022)Stewart, Robinson, Corley, Ortiz, Ferres, and Banerjee]{stewart2022torchgeo}
Adam~J Stewart, Caleb Robinson, Isaac~A Corley, Anthony Ortiz, Juan M~Lavista Ferres, and Arindam Banerjee.
\newblock Torchgeo: deep learning with geospatial data.
\newblock In \emph{Proceedings of the 30th international conference on advances in geographic information systems}, pp.\  1--12, 2022.

\bibitem[Sublime \& Kalinicheva(2019)Sublime and Kalinicheva]{sublime2019automatic}
J{\'e}r{\'e}mie Sublime and Ekaterina Kalinicheva.
\newblock Automatic post-disaster damage mapping using deep-learning techniques for change detection: Case study of the tohoku tsunami.
\newblock \emph{Remote Sensing}, 11\penalty0 (9):\penalty0 1123, 2019.

\bibitem[Tan \& Le(2019)Tan and Le]{tan2019efficientnet}
Mingxing Tan and Quoc Le.
\newblock Efficientnet: Rethinking model scaling for convolutional neural networks.
\newblock In \emph{International conference on machine learning}, pp.\  6105--6114. PMLR, 2019.

\bibitem[Tan et~al.(2024)Tan, Li, Chen, and Ai]{tan2024bd}
Yonghui Tan, Xiaolong Li, Yishu Chen, and Jinquan Ai.
\newblock Bd-msa: Body decouple vhr remote sensing image change detection method guided by multi-scale feature information aggregation.
\newblock \emph{arXiv preprint arXiv:2401.04330}, 2024.

\bibitem[Uy et~al.(2019)Uy, Pham, Hua, Nguyen, and Yeung]{uy2019revisiting}
Mikaela~Angelina Uy, Quang-Hieu Pham, Binh-Son Hua, Thanh Nguyen, and Sai-Kit Yeung.
\newblock Revisiting point cloud classification: A new benchmark dataset and classification model on real-world data.
\newblock In \emph{Proceedings of the IEEE/CVF international conference on computer vision}, pp.\  1588--1597, 2019.

\bibitem[Wan et~al.(2023)Wan, Tian, Kang, and Ma]{wan2023cldrnet}
Ling Wan, Ye~Tian, Wenchao Kang, and Lei Ma.
\newblock Cldrnet: A difference refinement network based on category context learning for remote sensing image change detection.
\newblock \emph{IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing}, 2023.

\bibitem[Wang et~al.(2023)Wang, Yang, Qiu, Liu, and Wu]{wang2023spatial}
Renfang Wang, Zijian Yang, Hong Qiu, Xiufeng Liu, and Dun Wu.
\newblock Spatial and channel exchange based on efficientnet for detecting changes of remote sensing images.
\newblock In \emph{2023 26th International Conference on Computer Supported Cooperative Work in Design (CSCWD)}, pp.\  1595--1600. IEEE, 2023.

\bibitem[Wang et~al.(2024)Wang, Wang, Hao, Wang, Wang, and Ye]{wang2024msgfnet}
Yukun Wang, Mengmeng Wang, Zhonghu Hao, Qiang Wang, Qianwen Wang, and Yuanxin Ye.
\newblock Msgfnet: Multi-scale gated fusion network for remote sensing image change detection.
\newblock \emph{Remote Sensing}, 16\penalty0 (3):\penalty0 572, 2024.

\bibitem[Watch(2002)]{watch2002global}
Global~Forest Watch.
\newblock Global forest watch.
\newblock \emph{World Resources Institute, Washington, DC Available from http://www. globalforestwatch. org (accessed March 2002)}, 2002.

\bibitem[Xu et~al.(2024)Xu, Zhang, Qi, Li, Yao, and Wang]{xu2024dual}
Zhongrong Xu, Chengkun Zhang, Jun Qi, Xilai Li, Bin Yao, and Lu~Wang.
\newblock A dual-difference change detection network for detecting building changes on high-resolution remote sensing images.
\newblock \emph{Geocarto International}, 39\penalty0 (1):\penalty0 2322080, 2024.

\bibitem[Yan et~al.(2024)Yan, Liu, Li, Zhang, and Yang]{yan2024multilevel}
Wenkai Yan, Yikun Liu, Mingsong Li, Ruifan Zhang, and Gongping Yang.
\newblock Multilevel feature aggregation and enhancement network for remote sensing change detection.
\newblock \emph{Journal of Applied Remote Sensing}, 18\penalty0 (1):\penalty0 016513--016513, 2024.

\bibitem[Yan et~al.(2023)Yan, Yue, Lin, Guo, Fang, Li, Xie, and Fang]{yan2023vectorization}
Yinglong Yan, Jun Yue, Jiaxing Lin, Zhengyang Guo, Yi~Fang, Zhenhao Li, Weiying Xie, and Leyuan Fang.
\newblock When vectorization meets change detection.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 2023.

\bibitem[Yang et~al.(2020)Yang, Xia, Liu, Du, Yang, Pelillo, and Zhang]{yang2020semantic}
Kunping Yang, Gui-Song Xia, Zicheng Liu, Bo~Du, Wen Yang, Marcello Pelillo, and Liangpei Zhang.
\newblock Semantic change detection with asymmetric siamese networks.
\newblock \emph{arXiv preprint arXiv:2010.05687}, 2020.

\bibitem[Yuan et~al.(2024)Yuan, Zhong, Yang, Li, and Dong]{yuan2024dynamically}
Shiying Yuan, Ruofei Zhong, Cankun Yang, Qingyang Li, and YaXin Dong.
\newblock Dynamically updated semi-supervised change detection network combining cross-supervision and screening algorithms.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 2024.

\bibitem[Zhang et~al.(2020)Zhang, Yue, Tapete, Jiang, Shangguan, Huang, and Liu]{zhang2020deeply}
Chenxiao Zhang, Peng Yue, Deodato Tapete, Liangcun Jiang, Boyi Shangguan, Li~Huang, and Guangchao Liu.
\newblock A deeply supervised image fusion network for change detection in high resolution bi-temporal remote sensing images.
\newblock \emph{ISPRS Journal of Photogrammetry and Remote Sensing}, 166:\penalty0 183--200, 2020.

\bibitem[Zhang et~al.(2024{\natexlab{a}})Zhang, Chen, Zhou, Chen, Liu, Zou, and Shi]{zhang2024bifa}
Haotian Zhang, Hao Chen, Chenyao Zhou, Keyan Chen, Chenyang Liu, Zhengxia Zou, and Zhenwei Shi.
\newblock Bifa: Remote sensing image change detection with bitemporal feature alignment.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 2024{\natexlab{a}}.

\bibitem[Zhang et~al.(2024{\natexlab{b}})Zhang, Li, Wang, and Wu]{zhang2024mfi}
Xingpeng Zhang, Yuru Li, Qiuli Wang, and Sijing Wu.
\newblock Mfi-cd: a lightweight siamese network with multidimensional feature interaction for change detection.
\newblock \emph{International Journal of Remote Sensing}, 45\penalty0 (8):\penalty0 2548--2566, 2024{\natexlab{b}}.

\bibitem[Zhao et~al.(2023{\natexlab{a}})Zhao, Zhao, Li, Song, and Wang]{zhao2023gampf}
Xiaoyang Zhao, Keyun Zhao, Siyao Li, Chuanming Song, and Xianghai Wang.
\newblock Gampf: A full-scale gated message passing framework based on collaborative estimation for vhr remote sensing image change detection.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 2023{\natexlab{a}}.

\bibitem[Zhao et~al.(2023{\natexlab{b}})Zhao, Zhao, Li, and Wang]{zhao2023gesanet}
Xiaoyang Zhao, Keyun Zhao, Siyao Li, and Xianghai Wang.
\newblock Gesanet: Geospatial-awareness network for vhr remote sensing image change detection.
\newblock \emph{IEEE Transactions on Geoscience and Remote Sensing}, 2023{\natexlab{b}}.

\bibitem[Zhao et~al.(2023{\natexlab{c}})Zhao, Zhang, Dong, and Du]{zhao2023adapting}
Yang Zhao, Yuxiang Zhang, Yanni Dong, and Bo~Du.
\newblock Adapting vision transformer for efficient change detection.
\newblock \emph{arXiv preprint arXiv:2312.04869}, 2023{\natexlab{c}}.

\bibitem[Zheng et~al.(2024)Zheng, Wu, Liu, Hung, and Wei]{zheng2024detail}
Dalong Zheng, Zebin Wu, Jia Liu, Chih-Cheng Hung, and Zhihui Wei.
\newblock Detail enhanced change detection in vhr images using a self-supervised multi-scale hybrid network.
\newblock \emph{IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing}, 2024.

\bibitem[Zhou et~al.(2024)Zhou, Qian, and Ren]{zhou2024multistage}
Meng Zhou, Weixian Qian, and Kan Ren.
\newblock Multistage interaction network for remote sensing change detection.
\newblock \emph{Remote Sensing}, 16\penalty0 (6):\penalty0 1077, 2024.

\end{thebibliography}

\documentclass{article} %
\usepackage{iclr2024_conference,times}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{hyperref}
\usepackage{url}
\usepackage{booktabs,graphicx,colortbl}

\definecolor{citecolor}{HTML}{0071BC}
\definecolor{linkcolor}{HTML}{ED1C24}
\definecolor{Gray}{gray}{0.86}

\hypersetup{colorlinks=True, citecolor=citecolor, urlcolor=magenta}

\title{A Change Detection Reality Check}

\iclrfinalcopy

\author{Isaac Corley\\
University of Texas at San Antonio\\
San Antonio, TX, USA \\
\texttt{isaac.corley@utsa.edu} \\
\And
Caleb Robinson \& Anthony Ortiz \\
Microsoft AI for Good Research Lab \\
Redmond, WA, USA \\
\texttt{\{caleb.robinson,anthony.ortiz\}@microsoft.com}
}

\begin{document}

\maketitle

\begin{abstract}
In recent years, there has been an explosion of proposed change detection deep learning architectures in the remote sensing literature. These approaches claim to offer state-of-the-art performance on different standard benchmark datasets. However, has the field truly made significant progress? In this paper we perform experiments which conclude a simple U-Net segmentation baseline without training tricks or complicated architectural changes is still a top performer for the task of change detection. All experiments are openly available at \href{https://github.com/isaaccorley/a-change-detection-reality-check}{github.com/isaaccorley/a-change-detection-reality-check}.
\end{abstract}

\section{Introduction}
\label{sec:intro}

The task of change detection from remotely sensed imagery is a canonical and important problem that allows us to analyze how our planet changes over time. It is crucial that change detection methods are highly accurate given their primary applications for building damage assessment disaster response~\citep{gupta2019creating,sublime2019automatic} and monitoring of the environment~\citep{cambrin2024cabuar,watch2002global}. The machine learning and remote sensing fields are currently experiencing their ImageNet~\citep{deng2009imagenet} benchmarking era with new proposed model architectures claiming incremental performance gains on standard benchmark datasets at a rapid rate. For example, the Changer paper~\citep{fang2023changer} -- first preprint released on Sep. 2022 -- is cited by 56 papers, the majority of them proposing a new architecture for change detection while also not releasing any open-source code or model checkpoints~\citep{ren2024mfinet,zhang2024mfi,zhou2024multistage,li4779358using,liu2024mutsimnet,huang2024remote,lin2024diformer,wang2024msgfnet,lu2024bi,tan2024bd,xu2024dual,yan2024multilevel,peng2023fda,wan2023cldrnet,zhao2023adapting,liu2023feature,quan2023unified,wang2023spatial,liu2023swin,fazry2023change}. Many of these proposed methods contain complicated architectural layers and modules specifically designed to better handle the bi-temporal image change detection task~\citep{bai2023deep,zhang2024bifa, yuan2024dynamically,li2024decoder,pang2024hicd,zheng2024detail,zhao2023gampf,li2023ms,yan2023vectorization,chen2023remote,liu2023attention,zhao2023gesanet}. Notably many of these papers are peer-reviewed and published at reputable conferences and journals. However, this begs the question if these new methods are actually improvements over generic segmentation architectures or have the benchmarks just been poorly executed? In this paper, we seek to answer this question.

\paragraph{Background}
When proposing a new model architecture it is common to perform a comparison to prior works through benchmark experiments to show statistically significant improvement. However over time, comparisons become less fair due to differences in training methodologies that are misconstrued as achieving state-of-the-art performance. Put simply, baseline experiments are often underpowered. This has been shown to be prevalent in many areas of research such as deep metric learning~\citep{musgrave2020metric}, unsupervised domain adaption~\citep{musgrave2021unsupervised}, image classification~\citep{bello2021revisiting}, deep reinforcement learning~\citep{henderson2018deep}, point cloud classification~\citep{uy2019revisiting}, and video recognition~\citep{du2021revisiting}. Furthermore, these analyses commonly conclude that simple baselines outperform complicated and task-specific architectures. Most recently, \citet{gerard2024simple} discovered that a generic architecture like U-Net, without training or evaluation tricks, is still competitive on the xBD dataset from the xView2 challenge~\citep{gupta2019xbd}. %

When analyzing relevant change detection papers and source code for benchmarking we find that it is common for authors to just compare to metrics reported in prior literature rather than re-running experiments with the prior methods in a consistent training setup. Further, we find that recent work often changes multiple aspects of the experimental setup, including the training routine (optimization methods, learning rate schedule, etc.) and loss functions, beyond just the proposed model architecture. This is problematic as any observed improvements on the benchmark dataset could be due to any combination of the new architecture, training routine, or loss function. This has the potential to result in unfair comparisons, especially if the improvement in quantitative performance is very small.

\textit{In this paper we revisit bi-temporal change detection benchmarks with simple baselines to get an overview on progress}. Specifically, we experiment with a simple semantic segmentation U-Net architecture, and siamese network variants~\citep{koch2015siamese}, to explore how these baselines perform against the latest state-of-the-art change detection methods. We find that this baseline architecture, from 2015, is a top performer on change detection benchmark datasets.

\section{Methods}
\label{methods}

\paragraph{State-of-the-Art Models}
In our experiments, in addition to compiling results from many prior change detection works, we retrain and evaluate the following state-of-the-art methods for change detection: BIT~\citep{chen2021remote} is a transformer-based siamese network architecture which uses a shared convolutional backbone to extract image features and transformer encoder decoder networks to perform change detection. ChangeFormer~\citep{bandara2022transformer} is an end-to-end transformer-based siamese network architecture for change detection. TinyCD~\citep{codegoni2023tinycd} is a change detection architecture which uses an EfficientNet~\citep{tan2019efficientnet} backbone to extract convolutional features to feed to a custom attention-based decoder network.

\paragraph{Our Baseline Models}

\citep{daudt2018fully} proposed three fully-convolutional (FC) architectures for change detection. The first is Early Fusion (FC-EF) which is an encoder-decoder style architecture with the change detection image pair concatenated as input. The other methods contain shared siamese encoders which either concatenate intermediate feature maps (FC-Siam-Conc) or take the difference (FC-Siam-Diff). While these networks are similar to the U-Net architecture~\citep{ronneberger2015u}, the original implementation is customized and hardcoded to be lightweight in parameter count and thus are unable to take advantage of pretrained encoder backbones like ResNet~\citep{he2016deep}. These methods are commonly seen in change detection benchmarks being compared to methods which utilize pretrained ImageNet backbones.

In our experiments, we generalize these implementations to use the standard U-Net framework which is able to take advantage of numerous ImageNet pretrained backbones. To avoid confusion with the original implementations of~\citep{daudt2018fully}, we refer to these architectures in our experiments as simply \texttt{U-Net}, \texttt{U-Net SiamConc}, and \texttt{U-Net SiamDiff}. Specifically, we utilize the implementations of these models in the TorchGeo~\citep{stewart2022torchgeo} library for reproducibility. We perform benchmarks using the ResNet-50 and EfficientNet-B4~\citep{tan2019efficientnet} backbones.

\begin{table}[t!]
\centering
\caption{Comparison of state-of-the-art and change detection architectures to a U-Net baseline on the LEVIR-CD dataset. We report the test set precision, recall, and F1 metrics of the positive change class. For the baseline experiments we perform 10 runs while varying random the seed and report metrics from the highest performing run. All other metrics are taken from their respective papers. The top performing methods are highlighted in bold. Gray rows indicate our baseline U-Net and siamese encoder variants.\\}
\resizebox{0.90\textwidth}{!}{%
\begin{tabular}{ccccc}
\toprule

\textbf{Model} &
\textbf{Backbone} &
\textbf{Precision} &
\textbf{Recall} &
\textbf{F1} \\
\toprule

FC-EF~\citep{daudt2018fully}         & -               & 86.91 & 80.17 & 83.40 \\
FC-Siam-Conc~\citep{daudt2018fully}   & -               & 91.99 & 76.77 & 83.69 \\
FC-Siam-Diff~\citep{daudt2018fully}   & -               & 89.53 & 83.31 & 86.31 \\
DTCDSCN~\citep{liu2020building}        & SE-Resnet34     & 88.53 & 86.83 & 87.67 \\
STANet~\citep{chen2020spatial}         & ResNet-18       & 83.81 & \textbf{91.00} & 87.26 \\
CDNet~\citep{chen2021adversarial}          & ResNet-18       & 91.60 & 86.50 & 89.00 \\
BIT~\citep{chen2021remote}            & ResNet-18       & 89.24 & 89.37 & 89.31 \\
ChangeFormer~\citep{bandara2022transformer}   & MiT-b1          & 92.59 & 89.68 & 91.11 \\
Tiny-CD~\citep{codegoni2023tinycd}        & EfficientNet-b4 & 92.68 & 89.47 & 91.05 \\
ChangerVanilla~\citep{fang2023changer} & ResNet-18       & 92.66 & 89.60 & 91.10 \\
ChangerEx~\citep{fang2023changer}      & ResNet-18       & 92.97 & 90.61 & \textbf{91.77} \\

\midrule
\midrule

\rowcolor{Gray}
\texttt{U-Net}~\citep{ronneberger2015u}          & EfficientNet-b4 & 92.69 & 87.16 & 89.25 \\
\rowcolor{Gray}
\texttt{U-Net}~\citep{ronneberger2015u}          & ResNet-50       & 91.97 & 89.78 & 90.38 \\
\rowcolor{Gray}
\texttt{U-Net SiamConc} & ResNet-50       & 92.87 & 89.48 & 90.41 \\
\rowcolor{Gray}
\texttt{U-Net SiamDiff} & ResNet-50       & \textbf{93.21} & 89.50 & 90.46 \\

\bottomrule
\end{tabular}%
}
\label{levircd-results}
\end{table}

\subsection{Datasets}
For comparisons of change detection architectures we utilize the following benchmark datasets:

\begin{description}
    \item[LEVIR-CD] \citep{chen2020spatial} A binary change detection dataset containing 637 high resolution (0.5m) 1024 × 1024 image pairs extracted from Google Earth. We utilize the splits provided with the dataset and, following other change detection papers, we convert the images to non-overlapping 256 × 256 patches.

    \item[WHU-CD] \citep{ji2018fully} A binary change detection dataset containing one pair of high-resolution (0.075m) aerial images of size 32507 × 15354. We utilize the train and test splits provided with the dataset and convert the images to non-overlapping 256 × 256 patches. During training runs we randomly split a 10\% holdout set from the train set to use as a validation set.
    
\end{description}

\section{Experiments}
\label{experiments}

\subsection{Baseline Training Details}
\label{sec:baseline-training-details}

Throughout our baseline training experiments we use the same hyperparameter setup from the BIT~\citep{chen2021remote} source which consists of batch size of 8, 200 epochs, stochastic gradient descent (SGD) optimizer with an initial learning rate $\gamma=0.01$, momentum $\mu=0.9$, weight decay $\alpha=5E-4$, and a linear decaying scheduler. We train each model by optimizing a cross entropy loss on a multiclass segmentation output where the number of classes is 2 for binary change detection. We select the checkpoint with the top performance based on the validation loss and do not perform any early stopping. We note that this setup results in a number of iterations which allows each model to converge.

During training we use augmentations consisting of random horizontal and vertical flips with probability $p=0.5$ and random resize crop with scale in the range $[0.8, 1.0]$, aspect ratio of $1$, and $p=1$. For image normalization we simply rescale the images to the range $[-1, 1]$ following the BIT~\citep{chen2021remote} source.

\subsection{Updated WHU-CD Benchmarks}
Many papers utilize a preprocessed and randomly split version of the WHU-CD dataset created in \citep{bandara2022revisiting} for change detection benchmarks. However these sets have been known to introduce data leakage~\citep{wgcbanchangeformergithubissue1} -- $\approx85\%$ of the test set is included in the train set due to a bug in the preprocessing scripts -- which makes them impossible to use to benchmark methods. We perform a new benchmark using the original train and test set splits from the WHU-CD dataset and retrain the BIT, ChangeFormer, and TinyCD models for comparison. We use the same experimental settings described in Section~\ref{sec:baseline-training-details}.

We find that performance of some models, particularly the transformer based methods BIT and ChangeFormer, can vary significantly over different random seeds, therefore we train each model on each dataset for runs with 10 different seeds. For comparisons we select the results from the highest performing run as well as report the average and standard deviation over runs for transparency. Any additional parameters can be found in our open source implementation.

\begin{table}[ht!]
\centering
\caption{Experimental results on the WHU-CD dataset. We retrain several state-of-the-art methods using the original dataset's train/test splits instead of the commonly used randomly split preprocessed version created in ~\cite{bandara2022revisiting}. We find that these state-of-the-art methods are outperformed by a U-Net baseline. We report the test set precision, recall, F1, and IoU metrics of the positive change class. For each run we select the model checkpoint with the lowest validation set loss. We provide metrics averaged over 10 runs with varying random seed as well as the best seed. Gray rows indicate our baseline U-Net and siamese encoder variants.\\}
\resizebox{0.90\textwidth}{!}{%
\begin{tabular}{cccccc}
\toprule
\textbf{Model} &
\textbf{Backbone} &
\textbf{F1} &
\textbf{Pre.} &
\textbf{Rec.} &
\textbf{IoU} \\
\toprule

\multicolumn{6}{l}{\textbf{Averaged Over 10 Seeds}} \\
\midrule
ChangeFormer & MiT-b1 & 75.65 ± 1.58 & 77.06 ± 3.22 & 74.67 ± 1.97 & 61.60 ± 2.05 \\
TinyCD & EfficientNet-b4 & 78.53 ± 1.28 & 80.15 ± 2.49 & 77.56 ± 2.13 & 65.52 ± 1.72 \\
BIT & ResNet-18 & 72.67 ± 2.69 & 70.30 ± 6.36 & 76.84 ± 4.53 & 58.06 ± 3.24 \\
\rowcolor{Gray}
\texttt{U-Net} & ResNet-50 & 81.85 ± 1.32 & 83.72 ± 2.65 & 80.39 ± 2.32 & 69.96 ± 1.83 \\
\rowcolor{Gray}
\texttt{U-Net SiamConc} & ResNet-50 & 81.33 ± 1.08 & 79.30 ± 2.78 & \textbf{84.19 ± 1.55} & 69.40 ± 1.52 \\
\rowcolor{Gray}
\texttt{U-Net SiamDiff} & ResNet-50 & \textbf{82.02 ± 1.48} & \textbf{83.82 ± 3.80} & 80.92 ± 2.51 & \textbf{70.29 ± 2.00} \\

\midrule
\multicolumn{6}{l}{\textbf{Best Seed}} \\
\midrule
ChangeFormer   & MiT-b1          & 77.75          & 82.60           & 78.57          & 64.22          \\
TinyCD         & EfficientNet-b4 & 78.53          & 80.15          & 77.56          & 65.52          \\
BIT            & ResNet-18       & 77.68          & 78.58          & 82.13          & 64.34          \\
\rowcolor{Gray}
\texttt{U-Net}          & ResNet-50       & \textbf{84.17} & \textbf{88.65} & 83.08          & \textbf{73.23} \\
\rowcolor{Gray}
\texttt{U-Net SiamConc} & ResNet-50       & 82.75          & 83.69          & \textbf{86.56} & 71.15          \\
\rowcolor{Gray}
\texttt{U-Net SiamDiff} & ResNet-50       & 84.01          & 88.56          & 85.63          & 73.02         \\

\bottomrule
\end{tabular}%
}
\label{whucd-results}
\end{table}

\section{Discussion}

\paragraph{Results}
In Table~\ref{levircd-results} we provide experimental results for our baseline on the LEVIR-CD dataset. We compare to prior work by compiling results from their respective papers. While the compiled results represent the best reported metrics for each previous method, it still is evident that a simple U-Net baseline is a top performer. Table~\ref{whucd-results} provides additional updated experimental results for the WHU-CD dataset and compares several state-of-the-art change detection architectures. Again, we can see that the baselines outperform all other methods.

Regarding the siamese variants of our U-Net baseline (\texttt{U-Net SiamDiff} and \texttt{U-Net SiamConc}), our experimental results indicate that processing each image in the bi-temporal pair with a shared encoder results in reasonable gains and is a promising approach. Methods such as Changer~\citep{fang2023changer} also conclude that additional feature interaction is key for achieving performance improvements for change detection.

\paragraph{Limitations}
Due to the fast pace of the field of change detection, with new architectures being proposed almost weekly, it is difficult to benchmark each new method. Furthermore, we do not compare to methods which propose a technique, e.g. a loss function, which is dependent on the architecture or methods, or to methods which do not publish open source code to reproduce the experimental results.

We note that there exists several other common change detection datasets such as DSIFN-CD~\citep{zhang2020deeply}, S2Looking~\citep{shen2021s2looking}, SECOND~\citep{yang2020semantic}, LEVIR-CD+~\citep{shen2021s2looking}, and xBD~\citep{gupta2019xbd}. While we do not benchmark against these, we leave this for future work.

\section{Conclusion and Next Steps}
In this paper we analyzed whether the field of change detection has actually made significant improvements on benchmark datasets in recent years. We conclude that many claimed improvements are questionable by demonstrating that a simple baseline of U-Net is still a top-performing method. To be clear, this is not an issue unique to change detection; other machine learning fields such as language modeling are finding it crucial to standardize fair benchmarking~\citep{srivastava2022beyond} when new methods are rapidly proposed. To mitigate this, we recommend utilizing and contributing proposed models to libraries and projects such as OpenCD~\citep{opencd}, GEO-Bench~\citep{lacoste2023geo}, and TorchGeo~\citep{stewart2022torchgeo} which standardize datasets and trainers for reliable benchmarking of remote sensing tasks. As this field has important downstream applications, we hope our results motivate the community to perform more reliable benchmarks of performance so that realistic advancements in change detection can be achieved. 

\subsubsection*{Acknowledgments}
We thank Jonathan Lwowski, Conor Wallace, Robin Cole, Sebastian Gerard, and Juan M. Lavista Ferres for their valuable feedback.

\bibliography{iclr2024_conference}
\bibliographystyle{iclr2024_conference}

\end{document}